OpenStudy (anonymous):

Factor: 121x^2 - 225y^4

OpenStudy (cwrw238):

this is difference of 2 squares apply the identity a^2 - b^2 = (a + b)(a - b) plug in a = sqrt 121x^2 and b = sqrt 225y^4